Calculate the pH of the resulting solution made by adding 2 moles of of KCN(aq) into 3 L of water.


1
Expert's answer
2021-01-20T03:17:23-0500

Concentration of KCN = 2 mol / 3 L .


In This Case Anionic hydrolysis So ,

PH = 7 + PK / 2 + log C / 2 .

= 7 + 9.2 / 2 + log 0.33

= 7 + 4.6 + ( - 0.4814 )

= 11.6 - 0.4814 = 11.6 - 0.5 = 11.1 Answer .